(Ignore income taxes in this problem.) Allen Company's required rate of return is 14%. The company is considering the purchase of a new machine that will save $10,000 per year in cash operating costs. The machine will cost $40,000 and will have an 8-year useful life with zero salvage value.


Required:

i) Compute the machine's internal rate of return to the nearest whole percent. Would you recommend purchase of the machine? Explain.


ii)     The company would like to use NPV to evaluate the project now. Compute the machine's NPV, assuming cost of capital is 10%. Would you recommend purchase of the machine? Explain.


iii) Discuss the implications for project investment priority based on your answer in (i) and (ii).





1
Expert's answer
2021-11-08T09:10:20-0500

Solution:

i.). Factor of the internal rate of return = Investment required "\div" Net annual cash flow

IRR factor = 40,000/10,000 = 4

Searching the 8th-period row in the PV annuity table, the value closest is 4.078 in the 18% column.

Actual IRR = 18.62"\%"

The machine should be purchased.

This is because the actual IRR is more than the required rate of return of 14"\%".

 

ii.). NPV:

Present value = 10,000 "\times" PV factor = 10,000 "\times" 5.3349 = 53,349

Less Initial Investment =                                                  (40,000)

NPV =                                                                                13,349

NPV = 13,349

The machine should be purchased.

This is because NPV is positive meaning that the discounted present value of the all-future cash flows associated with the investment will be positive and hence attractive.

 

iii.). Both methods signify that the investment is attractive and, therefore, Allen Company should purchase the new machine since it will generate positive returns.